OpenStudy (anonymous):

What are the factors of 64x2 + 48x + 9?

OpenStudy (aroub):

(8x+3)(8x+3)

OpenStudy (aroub):

= (8x+3)^2

OpenStudy (anonymous):

um can you show me the work so i can understand better please

OpenStudy (aroub):

There is MANY ways on how to factor. I'll show you my way. 64x2 + 48x + 9 What two number if you multiplied by each other you get 64x^2 ? 8x * 8x right ? so you put 8x in a bracket and the other 8x in a bracket like this : (8x+ ) (8x+ ) Now, what two number if you multiplied them by each other you get 9 ? 3*3 right ? so you put them in the brackets like this : (8x+3) (8x+3) which is (8x+3)^2 & for the signs they are all positive so the signs must be positive. How to check your answer ?  Multiply 8x by 3 and the other 8x by 3 8x*3= 24x 8x*3=24x Now add them 24x+24x = 48x Always when you add them you have to get the same number in the middle. Hope this helps :)